a. Calculate the mass of manganese (IV) oxide that can be synthesized from




15.00 grams of potassium iodide.




b. Calculate the percent yield of this experiment if a mass of 1.982 grams of




manganese (IV) oxide is produced.




c. Determine the percent error.




d. Compute for the excess amount of the excess reactant.

1
Expert's answer
2021-11-13T01:27:48-0500

A). 6KI(aq) +4H2O(l)+2KMnO4(aq)\to→ 3l2(s) +2 MnO2(s)+8KOH(aq)




REM of KI=166.00 g/mol


Moles of KI=15/ 166


A) = 0.09036moles

Mole ratio= KI:MnO2 = 6:2 = 3:1

Moles of MnO2= 0.09036/3

=0.03012moles

REM of MnO2= 86.99 g/mol

mass of MnO2= 0.03012× 86.97

=2.6195364g


B)Percent Yield =44.47%


C)Percent error is determined by the difference between the exact value and the approximate value of a quantity, divided by the exact value and then multiplied by 100 to represent it as a percentage of the exact value.


D)To find the amount of remaining excess reactant, subtract the mass of excess reagent consumed from the total mass of excess reagent given.